3x+2y-5z=1
-13y-24z=1
x+5y+8z=-5

Answer by addingup(3677)

3x+2y-5z=1
-13y-24z=1
x+5y+8z=-5
You answers:
x= 8 4/13
y= 5 6/13
z= -3
Sorry I'm out of time so I can't give you the step-by-step.